what's the point of intersection for the equations y=-4/3x-44/3 and y=-2x
Answer by stanbon(75887)

what's the point of intersection for the equations y=-4/3x-44/3 and y=-2x
-----------------
Substitute for "y" to get:
(-4/3)x- (44/3) = -2x
(-4/3)x - (44/3) = (-6/3)x
Add (6/3)x to both sides to get::
-------
(2/3)x = (44/3)
x = 22
---
Solve for "y":
y = -2x
y = -44
---------------
Cheers,
Stan H.
---------------